find a power series solution of xy'=y


1
Expert's answer
2022-02-15T18:18:35-0500

By Power Series Method, the solution of the differential equation is

"y=c_{0} \sum_{n=0}^{\infty} \frac{\left(\frac{x^{2}}{2}\right)^{n}}{n !}=c_{0} e^{\frac{x^{2}}{2}}," where "c_{0}" is any constant.

Let us look at some details.

"\text { Let } y=\sum_{n=0}^{\infty} c_{n} x^{n}"

By taking the derivative term by term,

"y^{\prime}=\sum_{n=1}^{\infty} n c_{n} x^{n-1}"

Now, let us look at the differential equation.

"y^{\prime}=x y"

by substituting the above power series in the equation,

"\Rightarrow \sum_{n=1}^{\infty} n c_{n} x^{n-1}=x \cdot \sum_{n=0}^{\infty} c_{n} x^{n}"  

by pulling the first term from the summation on the left,

"\Rightarrow c_{1}+\sum_{n=2}^{\infty} n c_{n} x^{n-1}=\sum_{n=0}^{\infty} c_{n} x^{n+1}"

by shifting the indices of the summation on the left by 2 ,

"\Rightarrow c_{1}+\sum_{n=0}^{\infty}(n+2) c_{n+2} x^{n+1}=\sum_{n=0}^{\infty} c_{n} x^{n+1}"

By matching coefficients,

"c_{1}=0" and "(n+2) c_{n+2}=c_{n} \Rightarrow c_{n+2}=\frac{c_{n}}{n+2}"

Let us observe the odd terms.

"\begin{aligned}

&c_{3}=\frac{c_{1}}{3}=\frac{0}{3}=0 \\

&c_{5}=\frac{c_{3}}{5}=\frac{0}{5}=0\\

&.\\

&.\\

&.\\

&c_{2 n+1}=0

\end{aligned}"

Let us observe the even terms.

"\begin{aligned}

&c_{2}=\frac{c_{0}}{2} \\

&c_{4}=\frac{c_{2}}{4}=\frac{c_{0}}{4 \cdot 2}=\frac{c_{0}}{2^{2} \cdot 2 !} \\

&c_{6}=\frac{c_{4}}{6}=\frac{c_{0}}{6 \cdot 4 \cdot 2}=\frac{c_{0}}{2^{3} \cdot 3 !}

&c_{2 n}=\frac{c_{0}}{2^{n} \cdot n !}

\end{aligned}"

Hence,

"y=\sum_{n=0}^{\infty} \frac{c_{0}}{2^{n} \cdot n !} x^{2 n}=c_{0} \sum_{n=0}^{\infty} \frac{\frac{x^{2 n}}{2^{n}}}{n !}=c_{0} \sum_{n=0}^{\infty} \frac{\left(\frac{x^{2}}{2}\right)^{n}}{n !}=c_{0} e^{\frac{x^{2}}{2}}"

"\text { (by replacing } x \text { by } \frac{x^{2}}{2} \text { in } e^{x}=\sum_{n=0}^{\infty} \frac{x^{n}}{n !} \text { ) }"